OpenStudy (anonymous):

Find the sum of the first 8 terms of the sequence. Show all work for full credit. 1, -3, -7, -11, ...

OpenStudy (anonymous):

Do you notice the pattern?

OpenStudy (anonymous):

Could you say what the next term is?

OpenStudy (anonymous):

here a=1, d= -3-1=-4, n=8 since\[s_n=\frac{n}{2}[2a+(n-1)d] \rightarrow s_8=\frac{8}{2}[2\times 1+(8-1)(-4)]\] \[=4[2+7(-4)] =4[2-28] =4[-26] =-104\]

OpenStudy (anonymous):

Seriously, bro?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Giving a student the answer almost never results in learning. So if you state the answer instead of leading them to find it on their own, your "teaching" is worthless.

OpenStudy (anonymous):

And in this case it nullifies others' efforts at actually teaching.

OpenStudy (ybarrap):

This is called an arithmetic series. The solution is easy to find when you see the pattern as @SmoothMath indicates. This will be very helpful: http://en.wikipedia.org/wiki/Arithmetic_series

OpenStudy (anonymous):

haha thanks guys, I appreciate the help @SmoothMath @dpasingh

OpenStudy (anonymous):

read this....it explains well http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm